what are the roots of the equation            x2-2x+3=0


pls solve so that i can understand the way of  solving this correctly............

Grade:10

2 Answers

Chetan Mandayam Nayakar
13 years ago

Dear Aditya, 

The roots of ax2+bx+c=0 are (-b±√(b2-4ac))/2a, tharefore the roots of x2-2x+3=0 are (2±√-8)/2 =1±(i√2)

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras

kirankumar kishor
13 years ago

you can use formula for this problem (-b +or- sqrt b^2-4ac)/2a. now we going to cal..

(-b=-2,a=1,c=3)

2=+or-sqrt 4(1)(3)/2(1)=2+sqrt-4/sqrt 2 or 2-sqrt-4/sqrt 2